### v1

@@ Abstract
In recent years, visually-rich document understanding has attracted increasing attention. Transformer-based pre-trained models have become the mainstream approach, yielding significant performance gains in this field. However, the self-attention mechanism’s quadratic computational complexity hinders their efficiency and ability to process long documents. In this paper, we present DocMamba, a novel framework based on the state space model. It is designed to reduce computational complexity to linear while preserving global modeling capabilities. To further enhance its effectiveness in document processing, we introduce the Segment-First Bidirectional Scan (SFBS) to capture contiguous semantic information. Experimental results demonstrate that DocMamba achieves new state-of-the-art results on downstream datasets such as FUNSD, CORD, and SORIE, while significantly improving speed and reducing memory usage. Notably, experiments on the HRDoc confirm DocMamba’s potential for length extrapolation. The code will be available online.
@@ Introduction
With the prosperity of commercial activities in today’s society,  a broad range of documents are used to convey information, leading to a growing demand for document processing (Cui et al. 2021). In order to reduce the labor-intensive workflows associated with this, Visually-rich Document Understanding (VrDU) (Xu et al. 2020a) is drawing considerable attention from both academia and industry. It aims to automate information extraction from documents and support various business applications.
@@
To achieve sub-quadratic complexity, one promising approach is the substitution of the Transformer with State Space Models (SSMs) (Gu et al. 2022a). They originated from the foundational classic state space model (Kalman 1960), and are notable for their capabilities in linear-time inference, highly parallelized training, and robust performance in tasks requiring long-context processing. Examples include the linear state space layers (LSSL) (Gu et al. 2021a), the structured state-space sequence model (S4) (Gu, Goel, and Ré 2021), and the diagonal state space (DSS) (Gupta, Gu, and Berant 2022). A recent addition to this category, Mamba (Gu and Dao 2023), has demonstrated exceptional results through its selective mechanism and hardware-aware design. Unlike self-attention mechanism in that each token interacts with all others within the context, Mamba enables each token to garner contextual knowledge solely through a compressed hidden state, thereby reducing the quadratic complexity to linear. Mamba has shown performance comparable to the Transformer in various fields, such as NLP (Lieber et al. 2024; Dat et al. 2024), image classification (Zhu et al. 2024; Liu et al. 2024), and speech processing (Zhang et al. 2024). Given the inherently longer sequences produced by documents, a natural question arises: Can Mamba work well for VrDU?
@@
We evaluate the performance of the pre-trained DocMamba using several publicly available benchmark datasets in downstream tasks. As depicted in Figure 1 (a), DocMamba surpasses the strong baseline LayoutLMv3 (Huang et al. 2022) in three datasets: the FUNSD dataset (Jaume, Ekenel, and Thiran 2019) for form understanding, the CORD (Park et al. 2019) dataset for receipt understanding, and the HRDoc (Ma et al. 2023) for semantic unit classification. Moreover, as shown in Figure 1 (b), tests on the HRDoc dataset show DocMamba has a faster inference speed and less GPU memory usage than LayoutLMv3. Especially with larger input lengths, DocMamba can save up to 88.3% of GPU memory and work 2.4 times faster, which can reduce the application costs significantly. This also proves the linear computational complexity of DocMamba. Furthermore, when the input length is restricted to 512 during both pre-training and fine-tuning, DocMamba still yields impressive results when the token length of test samples reaches 2560 for semantic unit classification on the HRDoc dataset. This validates DocMamba’s potential in length extrapolation. In conclusion, our research underscores the potential of SSMs as a powerful competitor with Transformer for VrDU, offering a simple yet effective baseline for future research.
@@
Early research (Yang et al. 2017; Katti et al. 2018) in VrDU typically utilizes unimodal or multimodal models with shallow fusion techniques. In recent years, the advent of pre-training techniques has revolutionized this field. BERT (Devlin et al. 2018) uses masked language models to obtain pre-trained deep bidirectional representations within pure text. Inspired by BERT, LayoutLM (Xu et al. 2020a) introduces 2-D spatial coordinate embeddings in addition to 1-D positional and text embeddings, thus simultaneously modeling the interaction between text and layout information within a singular framework. Furthermore, LayoutLMv2 (Xu et al. 2020b) adapts the standard Transformer (Vaswani et al. 2017) by integrating a spatial-aware self-attention mechanism, and concatenates visual tokens with textual tokens to enhance text-image interactions. LayoutLMv3 (Huang et al. 2022) suggests learning cross-modal alignment with unified text and image masking. Additionally, various model architectures (Appalaraju et al. 2021; Gu et al. 2022b), attention mechanisms (Hong et al. 2022; Zhang et al. 2022) and self-supervised tasks (Tu et al. 2023; Luo et al. 2023) have been explored, yielding promising results. However, nearly all of these methods are based on Transformer, which has a quadratic complexity concerning input length, thus posing challenges when processing lengthy documents.
@@
State Space Models (SSMs) serve as a fundamental model applied across various fields such as control theory (Raibert 1977), signal processing (Rao and Arun 1992), and applied economics (Schulz and Werwatz 2004). Recently, SSMs have garnered renewed attention within the deep learning community (Gu et al. 2021a; Gu, Goel, and Ré 2021; Smith, Warrington, and Linderman 2023), demonstrating notable proficiency in capturing long-range dependencies. They afford highly efficient computation, either as a recurrence or convolution operation, with linear or near-linear scalability in sequence length. Mamba (Gu and Dao 2023), in particular, distinguishes itself by incorporating a time-varying selection mechanism and a hardware-aware parallel algorithm. The significant potential demonstrated by Mamba has inspired a succession of studies in areas like NLP (Lieber et al. 2024; Dat et al. 2024), video understanding (Li et al. 2024; Lu, Salah, and Poppe 2024), speech processing (Zhang et al. 2024), and more. However, the application of SSMs for VrDU still remains unexplored.
@@
Table 1: Comparison with existing methods. “T/L/I” stands for “text/layout/image” modality. : UDoc split the CORD into 626/247 receipts for training/test, deviating from the official 800/100 split for training/test, so the score is not directly comparable. : TILT employed extra supervised data for pre-training, making its score not directly comparable as well.
@@
Comparison of Speed and Memory Usage. Among earlier Transformer-based methods, LayoutLMv3 stands out for its impressive performance and unified structure, making it our primary baseline method. To contrast the speed and memory consumption of DocMamba and LayoutLMv3, we choose HRDoc as the evaluation dataset for semantic unit classification. Figure 4 illustrates the memory consumption of both models during inference with the batch size set to 16. The memory consumption of LayoutLMv3 escalates rapidly, resulting in an Out-of-Memory situation when the input length reaches 3,072. Conversely, DocMamba’s memory consumption grows in a linear manner with the input length, saving 88.3% of memory when the input length attains 2,560. Figure 5 displays the inference speed of both models during inference. Batch size is set to 8 to avoid Out-of-Memory caused by LayoutLMv3. As the input length increases, the Frames Per Second (FPS) of LayoutLMv3 declines sharply. When the input length reaches 4096, DocMamba’s FPS becomes 2.4 times higher than that of LayoutLMv3. These results affirm the efficiency of DocMamba in processing text-dense documents, and also validate DocMamba’s linear computational complexity.
@@ Conclusion
In this study, we propose DocMamba, a model based on the SSM that does not rely on the self-attention mechanism. This reduces computational complexity to linear, making it suitable for processing text-dense documents. We also introduce Segment-First Bidirectional Scan, which is used to extract 1-D token sequences from documents. Experiments conducted on publicly available datasets, including FUNSD, CORD, and SROIE, show that DocMamba outperforms previous Transformer-based models, with faster speed and less memory usage. Further, outcomes on HRDoc validate DocMamba’s capacity for length extrapolation. This study highlights the potential of SSM as a powerful tool for understanding visually-rich documents and provides a simple yet effective baseline for future research.

### v2

@@ Abstract
In recent years, visually-rich document understanding has attracted increasing attention. Transformer-based pre-trained models have become the mainstream approach, yielding significant performance gains in this field. However, the self-attention mechanism’s quadratic computational complexity hinders their efficiency and ability to process long documents. In this paper, we present DocMamba, a novel framework based on the state space model. It is designed to reduce computational complexity to linear while preserving global modeling capabilities. To further enhance its effectiveness in document processing, we introduce the Segment-First Bidirectional Scan (SFBS) to capture contiguous semantic information. Experimental results demonstrate that DocMamba achieves new state-of-the-art results on downstream datasets such as FUNSD, CORD, and SORIE, while significantly improving speed and reducing memory usage. Notably, experiments on the HRDoc confirm DocMamba’s potential for length extrapolation.
Codehttps://github.com/Pengfei-Hu/DocMamba
@@ Introduction
With the prosperity of commercial activities in today’s society,  a broad range of documents are used to convey information, leading to a growing demand for document processing (Cui et al. 2021). In order to reduce the labor-intensive workflows associated with this, Visually-rich Document Understanding (VrDU) (Xu et al. 2020a) is drawing considerable attention from both academia and industry. It aims to automate information extraction from documents (Zhang et al. 2024b; Hu et al. 2024) and support various applications.
@@
To achieve sub-quadratic complexity, one promising approach is the substitution of the Transformer with State Space Models (SSMs) (Gu et al. 2022a). They originated from the foundational classic state space model (Kalman 1960), and are notable for their capabilities in linear-time inference, highly parallelized training, and robust performance in tasks requiring long-context processing. Examples include the linear state space layers (LSSL) (Gu et al. 2021a) and the structured state-space sequence model (S4) (Gu, Goel, and Ré 2021). A recent addition to this category, Mamba (Gu and Dao 2023), has demonstrated exceptional results through its selective mechanism and hardware-aware design. Unlike self-attention mechanism in that each token interacts with all others within the context, Mamba enables each token to garner contextual knowledge solely through a compressed hidden state, thereby reducing the quadratic complexity to linear. Mamba has shown performance comparable to the Transformer in various fields (Lieber et al. 2024; Zhu et al. 2024). Given the inherently longer sequences produced by documents, a natural question arises: Can Mamba work well for VrDU?
@@
We evaluate the performance of the pre-trained DocMamba using several publicly available benchmark datasets in downstream tasks. As depicted in Figure 1 (a), DocMamba surpasses the strong baseline LayoutLMv3 (Huang et al. 2022) at the base scale with a similar number of parameters across three datasets: the FUNSD dataset (Jaume, Ekenel, and Thiran 2019) for form understanding, the CORD (Park et al. 2019) dataset for receipt understanding, and the HRDoc (Ma et al. 2023) for semantic unit classification. Moreover, as shown in Figure 1 (b), tests on the HRDoc dataset show DocMamba has a faster inference speed and less GPU memory usage than LayoutLMv3. Especially with larger input lengths, DocMamba can save up to 88.3% of GPU memory and work 2.4 times faster, which can reduce the application costs significantly. This also proves the linear computational complexity of DocMamba. Furthermore, when the input length is restricted to 512 during both pre-training and fine-tuning, DocMamba still yields impressive results when the token length of test samples reaches 2560 for semantic unit classification on the HRDoc dataset. This validates DocMamba’s potential in length extrapolation. In conclusion, our research underscores the potential of SSMs as a powerful competitor with Transformer for VrDU, offering a simple yet effective baseline for future research.
@@
Early research (Yang et al. 2017; Hu et al. 2022) in VrDU typically utilizes unimodal or multimodal models with shallow fusion techniques. In recent years, the advent of pre-training techniques has revolutionized this field. BERT (Devlin et al. 2018) uses masked language models to obtain pre-trained deep bidirectional representations within pure text. Inspired by BERT, LayoutLM (Xu et al. 2020a) introduces 2-D spatial coordinate embeddings in addition to 1-D positional and text embeddings, thus simultaneously modeling the interaction between text and layout information within a singular framework. Furthermore, LayoutLMv2 (Xu et al. 2020b) adapts the standard Transformer by integrating a spatial-aware self-attention mechanism, and concatenates visual tokens with textual tokens to enhance text-image interactions. LayoutLMv3 (Huang et al. 2022) suggests learning cross-modal alignment with unified text and image masking. Additionally, various model architectures (Appalaraju et al. 2021; Gu et al. 2022b), attention mechanisms (Hong et al. 2022; Zhang et al. 2022) and self-supervised tasks (Tu et al. 2023; Luo et al. 2023; Zhang et al. 2024c; Yao, Li, and Xiao 2024) have been explored. However, nearly all of these methods are based on Transformer, which has a quadratic complexity concerning input length, thus posing challenges when processing lengthy documents.
@@
State Space Models (SSMs) serve as a fundamental model applied across various fields such as control theory (Raibert 1977), signal processing (Rao and Arun 1992), and applied economics (Schulz and Werwatz 2004). Recently, SSMs have garnered renewed attention within the deep learning community (Gu et al. 2021a; Gu, Goel, and Ré 2021; Smith, Warrington, and Linderman 2023), demonstrating notable proficiency in capturing long-range dependencies. They afford highly efficient computation, either as a recurrence or convolution operation, with linear or near-linear scalability in sequence length. Mamba (Gu and Dao 2023), in particular, distinguishes itself by incorporating a time-varying selection mechanism and a hardware-aware parallel algorithm. The significant potential demonstrated by Mamba has inspired a succession of studies in areas like NLP (Lieber et al. 2024; Dat et al. 2024), video understanding (Li et al. 2024; Lu, Salah, and Poppe 2024; Yao et al. 2024), speech processing (Zhang et al. 2024a), and more. However, the application of SSMs for VrDU still remains unexplored.
@@
Table 1: Comparison with existing methods. “T/L/I” stands for “text/layout/image” modality. : UDoc split the CORD into 626/247 receipts for training/test, deviating from the official 800/100 split for training/test, so the score is not directly comparable. : TILT employed extra supervised data for pre-training, making its score not directly comparable as well. : To keep a fair comparison with DocMamba, we use the same data as DocMamba to pretrain the vanilla Mamba from scratch.
@@
Comparison of Speed and Memory Usage. Among earlier Transformer-based methods, LayoutLMv3 stands out for its impressive performance and unified structure, making it our primary baseline method. To contrast the speed and memory consumption of DocMamba and LayoutLMv3, we choose HRDoc as the evaluation dataset for semantic unit classification. We use the official implementation of LayoutLMv3 available on Hugging Face  [^3] https://huggingface.co/microsoft/layoutlmv3-base for benchmarking. Figure 4 illustrates the memory consumption of both models during inference with the batch size set to 16. The memory consumption of LayoutLMv3 escalates rapidly, resulting in an Out-of-Memory situation when the input length reaches 3,072. Conversely, DocMamba’s memory consumption grows in a linear manner with the input length, saving 88.3% of memory when the input length attains 2,560. Figure 5 displays the inference speed of both models during inference. Batch size is set to 8 to avoid Out-of-Memory caused by LayoutLMv3. As the input length increases, the Frames Per Second (FPS) of LayoutLMv3 declines sharply. When the input length reaches 4096, DocMamba’s FPS becomes 2.4 times higher than that of LayoutLMv3. These results affirm the efficiency of DocMamba in processing text-dense documents, and also validate DocMamba’s linear computational complexity.
@@ Conclusion
In this study, we propose DocMamba, a model based on the SSM that does not rely on the self-attention mechanism. This reduces computational complexity to linear, making it suitable for processing text-dense documents. We also introduce Segment-First Bidirectional Scan, which is used to extract 1-D token sequences from documents. In addition, DocMamba combines text and layout information using a multi-layer bidirectional Mamba encoder. Experiments conducted on publicly available datasets, including FUNSD, CORD, and SROIE, show that DocMamba outperforms previous Transformer-based models, with faster speed and less memory usage. Further, outcomes on HRDoc validate DocMamba’s capacity for length extrapolation. This study highlights the potential of SSM as a powerful tool for understanding visually-rich documents and provides a simple yet effective baseline for future research.